Solve the following problem showing your work.
Geometry. If the sides of a square are decreased by 2 cm, the area is decreased by 36 cm^2. What were the dimensions of the original square?

Answer by ankor@dixie-net.com(22740)

solve the following problem showing your work.
Geometry. If the sides of a square are decreased by 2 cm, the area is decreased by 36 cm^2. What were the dimensions of the original square?
:
Let x = side of the original square:
then, x^2 = area of the original square
:
(x-2) = side of the reduced square;
then (x-2)^2 = area of the reduced square
:
problems says:
Original area - reduced area = 36 sq cm
x^2 - (x-2)^2 = 36
:
FOIL (x-2)(x-2)
x^2 - (x^2 - 4x + 4) = 36
:
Remove brackets, (change the signs inside the brackets)
x^2 - x^2 + 4x - 4 = 36
:
Fortunately the x^2's eliminate themselves!
4x = 36 + 4
4x = 40
x = 40/4
x = 10 cm is the dimension of the original square
:
Can easily be checked:
10^2 - 8^2 =
100 - 64 = 36
:
How about this? Could you follow this OK? Any questions?